# Mind the Gap: Bridging Thought Leap for Improved Chain-of-Thought Tuning

### v2

@@ Abstract
Large language models (LLMs) have achieved remarkable progress on mathematical tasks through Chain-of-Thought (CoT) reasoning. However, existing mathematical CoT datasets often suffer from Thought Leaps due to experts omitting intermediate steps, which negatively impacts model learning and generalization. We propose the CoT Thought Leap Bridge Task, which aims to automatically detect leaps and generate missing intermediate reasoning steps to restore the completeness and coherence of CoT. To facilitate this, we constructed a specialized training dataset called ScaleQM+, based on the structured ScaleQuestMath dataset, and trained CoT-Bridge to bridge thought leaps. Through comprehensive experiments on mathematical reasoning benchmarks, we demonstrate that models fine-tuned on bridged datasets consistently outperform those trained on original datasets, with improvements of up to +5.87% on NuminaMath. Our approach effectively enhances distilled data (+3.02%) and provides better starting points for reinforcement learning (+3.1%), functioning as a plug-and-play module compatible with existing optimization techniques. Furthermore, CoT-Bridge demonstrate improved generalization to out-of-domain logical reasoning tasks, confirming that enhancing reasoning completeness yields broadly applicable benefits.
@@ 1 Introduction
Large Language Models (LLMs) [1, 2, 3] have demonstrated significant performance improvements through Chain-of-Thought (CoT) reasoning [4], particularly on complex tasks such as mathematics and coding [5]. CoT guides models to generate intermediate reasoning steps, simulating the human process of step-by-step problem solving, thereby enhancing both solution accuracy and interpretability [6]. The quality of these reasoning chains directly impacts model performance, serving as a critical foundation for advanced reasoning capabilities.
@@ 2.2 Thought Leap Bridging Dataset Construction and Bridge Model Training
To implement the CoT Thought Leap Bridge task, we created a specialized training dataset named ScaleQM+ based on the structurally comprehensive ScaleQuestMath dataset. As illustrated in Figure x, we start with complete reasoning chains $C^{*}=(s_{0},s_{1}^{*},\dots,s_{m}^{*})$ where $s_{0}=Q$ is the question, and strategically remove intermediate steps to produce incomplete chains $C=(s_{0},s_{1},\dots,s_{n})$ containing Thought Leaps. The removed steps form the reference set of missing intermediate steps needed to bridge these leaps.
@@
For each modified chain $C$, we generate two essential training components: the ground-truth Thought Leap positions $\mathcal{L}_{gt}$ identifying where step deletion resulted in reasoning gaps, and the corresponding ground-truth missing steps $\mathcal{M}_{gt}$ consisting of the original removed steps. Our primary model, CoT-Bridge, learns the mapping $f:C\rightarrow(\hat{\mathcal{L}},\hat{\mathcal{M}})$, taking an incomplete chain as input and producing both predicted leap positions and the corresponding missing steps. As a baseline for comparison, we also implement CoT-Bridge-Random, which learns $g:(C,\mathcal{L}_{gt})\rightarrow\hat{\mathcal{M}}$, focusing solely on generating missing steps given the ground-truth leap positions.
@@ 2.3 Data Augmentation with Bridge Model
After training, we apply CoT-Bridge to enhance existing mathematical reasoning datasets, specifically MetaMathQA and NuminaMath-CoT, creating improved versions named MetaMath-Bridge and NuminaMath-Bridge respectively. As shown in Figure x, CoT-Bridge processes reasoning chains to identify gaps and generate appropriate bridging content.
@@ 3.3 Baselines
We establish a series of baselines for comparison. The most important baseline is standard SFT performed directly on the original datasets. In addition, we use a zero-shot briding which method leveraging general-purpose LLMs (in this paper, Qwen2.5-Instruct-7B/72B [23]) to explore Thought Leap bridging without direct training. This baseline aims to assess the effectiveness of pure prompt engineering in bridging Thought Leaps, determining whether specialized detection and repair mechanisms are necessary. Furthermore, we use CoT-Bridge-Random to generated intermediate steps at randomly chosen positions, to evaluate the importance of accurate Thought Leap localization for effective bridging. As a reference, we consider the performance of base models in a 4-shot setting. Additionally, we present the results of SFT using GSM8K+MATH and MathInstruct.
@@ 4.2 Evaluation of Thought Leap Bridge Task
To evaluate the capability of various methods in CoT Thought Leap Bridge Task, we constructed a standardized evaluation framework on ScaleQM+ test set, covering leap identification and generation quality. Calculation methods for these metrics are detailed in Appendix H. As shown in Table 3 CoT-Bridge significantly outperforms other methods in localization precision (78.02%), redundancy rate (1.61%) and overall metric (76.15%). Qwen2.5-Instruct-72B notably outperforms the 7B model, indicating that stronger generative models possess certain zero-shot localization and restoration capabilities. We also evaluated an exhaustive approach based on full-position generation and similarity filtering. This method uses CoT-Bridge-Random to generate candidate content between all adjacent steps, filtering them using similarity thresholds. While this approach demonstrates high recall capability, it exhibits an extremely high redundancy rate. The substantial redundant content may cause LLMs to learn low-quality repetitive patterns during fine-tuning, thereby compromising reasoning ability. Additionally, this strategy is highly sensitive to similarity thresholds, incurs significant computational costs, and is difficult to scale in practice.
@@ 4.3 Evaluation of Out-of-Domain Reasoning Capability
As show in Table 4, we compared the performance of models trained on NuminaMath and NuminaMath-Bridge. Results indicate that CoT-Bridge improved OOD reasoning performance (Meta-Llama3.1-8B: $\uparrow$2.99%, Qwen2.5-Math-1.5B: $\uparrow$0.99%) while reducing invalid response rates. We attribute the performance improvement to the refined intermediate steps introduced by CoT-Bridge, which help models master general reasoning structures and enhance generalization capabilities.
@@ 4.4 Analysis of Bridging Positions
Statistics show that middle type bridgings dominate (MetaMathQA: 66.34%; NuminaMath: 55.52%), followed by begin, with end accounting for a relatively small proportion (approximately 2%). To evaluate the specific role of each type of content, we designed position component ablation experiments: removing one type of bridging (begin / middle / end) from the CoT-Bridge data while keeping the rest unchanged, and fine-tuning Qwen2.5-Math-1.5B under identical training settings. As show in Table 5, experimental results indicate that removal leads to performance degradation. This suggests that all three types of completion have a positive effect on model performance. We believe that begin bridgings help clarify problem objectives and establish solution frameworks; middle bridgings support critical links in reasoning paths; and end bridgings, although proportionally smaller, help strengthen logical closure and result reasonability in certain tasks.
@@ 4.5 Process Supervision Scoring and Noise Impact Analysis
To measure the bridging quality of various methods, we introduced the process supervision model Qwen2.5-Math-PRM-7B to score the intermediate steps (range 0–1) and analyzed the score distribution. Table 11 in Appendix J and Figure 4 present the PRM score distribution of the bridged steps. Results show that steps generated by CoT-Bridge have higher quality compared to other methods. On the NuminaMath dataset, CoT-Bridge’s proportion in the high-quality (range 0.9–1) reaches 83.51%, significantly higher than 72B Fill (53.45%) and 7B Fill (56.8%). Furthermore, CoT-Bridge’s proportion in the low-quality (range 0–0.1) is only 2.08%, compared to 14.81% and 13.12% for 72B and 7B Fill, respectively. We also used DeepSeek-R1 [3] to score the CoT comprehensively, CoT bridged by CoT-Bridge scoring higher than the original chains. The scoring prompt, results and examples are provided in the Appendix I.
@@ 5 Conclusion
In this paper, we addressed the critical issue of Thought Leaps in CoT reasoning by introducing the CoT Thought Leap Bridge Task and developing the CoT-Bridge model trained on ScaleQM+ dataset. Our approach automatically detects and fills these reasoning gaps, demonstrably enhancing the completeness and coherence of reasoning chains. Comprehensive experiments revealed that models tuned on these bridged datasets leads to significant performance improvements and better generalization to out-of-domain reasoning tasks. Furthermore, CoT-Bridge acts as an effective plug-and-play module, improving outcomes in knowledge distillation and reinforcement learning, thereby highlighting the substantial benefits of rectifying thought leaps for more robust and capable large language models.
@@ A.1 Methods for Enhancing Mathematical Reasoning
Inference-time approaches improve reasoning by adopting advanced decoding strategies or utilizing extra computational resources. Self-Consistency (SC)[32] enhances robustness by sampling multiple reasoning paths and applying majority voting. Search-based methods, including Tree-of-Thought (ToT)[33] and Monte Carlo Tree Search (MCTS)[34], systematically explore reasoning spaces for optimal solutions. Recent test-time expansion techniques, such as multi-round thinking[35] and s1 budget forcing[36], dynamically increase computation to boost performance.
Training-based approaches internalize enhanced reasoning abilities through parameter updates during model training. A common strategy involves fine-tuning models on datasets with explicitly detailed intermediate reasoning steps. Additionally, training paradigms incorporating self-reflection, self-verification, and self-correction have gained attention, teaching models to recognize and rectify errors autonomously. Frameworks like StepCo[37], S2R[38], and ReVISE[39] integrate these metacognitive capabilities through supervised, reinforcement, or preference learning. We leverage CoT-Bridge to enhance the structural completeness of existing datasets, enabling models to acquire finer-grained reasoning patterns during training.
@@ D.2 Evaluation Settings
The maximum token limit for generation was set to 2048, and prompt templates remained consistent with training ones. It is noteworthy that, although we set the decoding temperature parameter to zero, vLLM outputs still exhibited some randomness. For answer extraction and comparison, we employed the Math-Verify[47] tool. Given limitations of math-verify in handling complex expressions, responses failing initial verification were subsequently validated using DeepSeek R1[3]. All model evaluations were performed using 4 NVIDIA A100-40G GPUs.
@@ G.2 Results
We present the detailed results from Figure 4(b) in Table 8. For simplicity, only Math-Verify is used for answer verification without further applying DeepSeek R1.
@@ H.2 Overall Metric
Appendix I CoT Score
I.1 Score Template
Since scoring a single response with DeepSeek R1 can result in high variance, we instead compare the original CoT and its bridged version for the same question side-by-side.
DeepSeek R1 Score Template You are a math reasoning expert. Please evaluate the coherence of the given two Chain of Thoughts (CoTs) based on the provided question.
@@ I.3 Score Result
Table 10: DeepSeek R1 scoring results for the original CoT and the bridged CoT.
Appendix J PRM Score

### v3

@@ Abstract
Large language models (LLMs) have achieved remarkable progress on mathematical tasks through Chain-of-Thought (CoT) reasoning. However, existing mathematical CoT datasets often suffer from Thought Leaps due to experts omitting intermediate steps, which negatively impacts model learning and generalization. We propose the CoT Thought Leap Bridge Task, which aims to automatically detect leaps and generate missing intermediate reasoning steps to restore the completeness and coherence of CoT. To facilitate this, we constructed a specialized training dataset called ScaleQM+, based on the structured ScaleQuestMath dataset, and trained CoT-Bridge to bridge thought leaps. Through comprehensive experiments on mathematical reasoning benchmarks, we demonstrate that models fine-tuned on bridged datasets consistently outperform those trained on original datasets, with improvements of up to +5.87% on NuminaMath. Our approach effectively enhances distilled data (+3.02%) and provides better starting points for reinforcement learning (+3.1%), functioning as a plug-and-play module compatible with existing optimization techniques. Furthermore, CoT-Bridge demonstrates improved generalization to out-of-domain logical reasoning tasks, confirming that enhancing reasoning completeness yields broadly applicable benefits.
@@ 1 Introduction
Large Language Models (LLMs) [1, 2, 3, 4, 5, 6] have demonstrated significant performance improvements through Chain-of-Thought (CoT) reasoning [7, 8], particularly on complex tasks [9] such as mathematics [10] and coding. CoT guides models to generate intermediate reasoning steps, simulating the human process of step-by-step problem solving, thereby enhancing both solution accuracy and interpretability [11]. The quality of these reasoning chains directly impacts model performance, serving as a critical foundation for advanced reasoning capabilities.
@@ 2.2 Thought Leap Bridging Dataset Construction and Bridge Model Training
To implement the CoT Thought Leap Bridge task, we created a specialized training dataset named ScaleQM+ based on the structurally comprehensive ScaleQuestMath dataset. We chose ScaleQuestMath as an approximation of ideal CoT because its reasoning structure is relatively complete and well-formed, making it suitable for systematic step removal. As illustrated in Figure 2, we start with complete reasoning chains $C^{*}=(s_{0},s_{1}^{*},\dots,s_{m}^{*})$ where $s_{0}=Q$ is the question, and strategically remove intermediate steps to produce incomplete chains $C=(s_{0},s_{1},\dots,s_{n})$ containing Thought Leaps. The removed steps form the reference set of missing intermediate steps needed to bridge these leaps.
@@
[…]
@@ 2.3 Data Augmentation with Bridge Model
After training, we apply CoT-Bridge to enhance existing mathematical reasoning datasets, specifically MetaMathQA and NuminaMath-CoT, creating improved versions named MetaMath-Bridge and NuminaMath-Bridge respectively. As shown in Figure 2, CoT-Bridge processes reasoning chains to identify gaps and generate appropriate bridging content.
@@ 3.3 Baselines
We establish a series of baselines for comparison. The most important baseline is standard SFT performed directly on the original datasets. In addition, we use a zero-shot bridging which method leveraging general-purpose LLMs (in this paper, Qwen2.5-Instruct-7B/72B [28]) to explore Thought Leap bridging without direct training. This baseline aims to assess the effectiveness of pure prompt engineering in bridging Thought Leaps, determining whether specialized detection and repair mechanisms are necessary. Furthermore, we use CoT-Bridge-Random to generated intermediate steps at randomly chosen positions, to evaluate the importance of accurate Thought Leap localization for effective bridging. As a reference, we consider the performance of base models in a 4-shot setting. Additionally, we present the results of SFT using GSM8K+MATH and MathInstruct.
@@ 4.2 Evaluation of Thought Leap Bridge Task
To evaluate the capability of various methods in CoT Thought Leap Bridge Task, we constructed a standardized evaluation framework on ScaleQM+ test set, covering leap identification and generation quality. Calculation methods for these metrics are detailed in Appendix H. Case study is detailed in Appendix I. As shown in Table 3, CoT-Bridge significantly outperforms other methods in localization precision (78.02%), redundancy rate (1.61%) and overall metric (76.15%). Qwen2.5-Instruct-72B notably outperforms the 7B model, indicating that stronger generative models possess certain zero-shot localization and restoration capabilities. We also evaluated an exhaustive approach based on full-position generation and similarity filtering. This method uses CoT-Bridge-Random to generate candidate content between all adjacent steps, filtering them using similarity thresholds. While this approach demonstrates high recall capability, it exhibits an extremely high redundancy rate. The substantial redundant content may cause LLMs to learn low-quality repetitive patterns during fine-tuning, thereby compromising reasoning ability. Additionally, this strategy is highly sensitive to similarity thresholds, incurs significant computational costs, and is difficult to scale in practice.
@@ 4.3 Evaluation of Out-of-Domain Reasoning Capability
As shown in Table 4, we compared the performance of models trained on NuminaMath and NuminaMath-Bridge. Results indicate that CoT-Bridge improved OOD reasoning performance (Meta-Llama3.1-8B: $\uparrow$2.99%, Qwen2.5-Math-1.5B: $\uparrow$0.99%) while reducing invalid response rates. We attribute the performance improvement to the refined intermediate steps introduced by CoT-Bridge, which help models master general reasoning structures and enhance generalization capabilities.
@@ 4.4 Analysis of Bridging Positions
Statistics show that middle type bridgings dominate (MetaMathQA: 66.34%; NuminaMath: 55.52%), followed by begin, with end accounting for a relatively small proportion (approximately 2%). To evaluate the specific role of each type of content, we designed position component ablation experiments: removing one type of bridging (begin / middle / end) from the CoT-Bridge data while keeping the rest unchanged, and fine-tuning Qwen2.5-Math-1.5B under identical training settings. As shown in Table 5, experimental results indicate that removal leads to performance degradation. This suggests that all three types of completion have a positive effect on model performance. We believe that begin bridgings help clarify problem objectives and establish solution frameworks; middle bridgings support critical links in reasoning paths; and end bridgings, although proportionally smaller, help strengthen logical closure and result reasonability in certain tasks.
@@ 4.5 Process Supervision Scoring and Noise Impact Analysis
To measure the bridging quality of various methods, we introduced the process supervision model Qwen2.5-Math-PRM-7B to score the intermediate steps (range 0–1) and analyzed the score distribution. Table 11 in Appendix K and Figure 4 present the PRM score distribution of the bridged steps. Results show that steps generated by CoT-Bridge have higher quality compared to other methods. On the NuminaMath dataset, CoT-Bridge’s proportion in the high-quality (range 0.9–1) reaches 83.51%, significantly higher than 72B Fill (53.45%) and 7B Fill (56.8%). Furthermore, CoT-Bridge’s proportion in the low-quality (range 0–0.1) is only 2.08%, compared to 14.81% and 13.12% for 72B and 7B Fill, respectively. We also used DeepSeek-R1 [6] to score the CoT comprehensively, CoT bridged by CoT-Bridge scoring higher than the original chains. The scoring prompt, results and examples are provided in the Appendix J.
@@ 5 Conclusion
In this paper, we addressed the critical issue of Thought Leaps in CoT reasoning by introducing the CoT Thought Leap Bridge Task and developing the CoT-Bridge model trained on ScaleQM+ dataset. Our approach automatically detects and fills these reasoning gaps, demonstrably enhancing the completeness and coherence of reasoning chains. Comprehensive experiments revealed that models tuned on these bridged datasets lead to significant performance improvements and better generalization to out-of-domain reasoning tasks. Furthermore, CoT-Bridge acts as an effective plug-and-play module, improving outcomes in knowledge distillation and reinforcement learning, thereby highlighting the substantial benefits of rectifying thought leaps for more robust and capable large language models.
Acknowledgement
This work is supported by the National Natural Science Foundation of China (No. 62436007), the Key Research and Development Program of Zhejiang Province, China (No. 2024C01034), the Fundamental Research Funds for the Central Universities (226-2024-00170), MOE Engineering Research Center of Digital Library, CCF-Baidu Open Fund and ZJU Kunpeng&Ascend Center of Excellence.
@@ A.1 Methods for Enhancing Mathematical Reasoning
Inference-time approaches improve reasoning by adopting advanced decoding strategies or utilizing extra computational resources. Self-Consistency (SC) [37] enhances robustness by sampling multiple reasoning paths and applying majority voting. Search-based methods, including Tree-of-Thought (ToT) [38] and Monte Carlo Tree Search (MCTS) [39], systematically explore reasoning spaces for optimal solutions. Recent test-time expansion techniques, such as multi-round thinking [40] and s1 budget forcing [41], dynamically increase computation to boost performance. In addition, inference-time intervention methods [42, 43] enhance reasoning performance by manipulating hidden states.
Training-based approaches internalize enhanced reasoning abilities through parameter updates during model training. A common strategy involves fine-tuning models on datasets with explicitly detailed intermediate reasoning steps. Additionally, training paradigms incorporating self-reflection [44], self-verification, self-breaking [45] and self-correction have gained attention, teaching models to recognize and rectify errors autonomously. Frameworks like StepCo [46], S2R [47], and ReVISE [48] integrate these metacognitive capabilities through supervised, reinforcement, or preference learning. In particular, recent advances highlight that reinforcement learning plays an especially crucial role [49, 50, 51, 52] in enhancing the overall capabilities of large models. We leverage CoT-Bridge to enhance the structural completeness of existing datasets, enabling models to acquire finer-grained reasoning patterns during training and to be seamlessly integrated into existing training pipelines.
@@ D.2 Evaluation Settings
The maximum token limit for generation was set to 2048, and prompt templates remained consistent with training ones. It is noteworthy that, although we set the decoding temperature parameter to zero, vLLM outputs still exhibited some randomness. For answer extraction and comparison, we employed the Math-Verify [60] tool. Given limitations of math-verify in handling complex expressions, responses failing initial verification were subsequently validated using DeepSeek-R1 [6]. All model evaluations were performed using 4 NVIDIA A100-40G GPUs.
@@ G.2 Results
We present the detailed results from Figure 4(b) in Table 8. For simplicity, only Math-Verify is used for answer verification without further applying DeepSeek-R1.
@@ H.2 Overall Metric
Appendix I Case Study: CoT Thought Leap Bridge
Following the example in Figure 1(a), we compared the zero-shot bridging results of Qwen2.5-Instruct-7B/72B, GPT-4o, and Gemini-2.5-PRO.
CoT-Bridge Missing Step 1:
The missing step should be placed between Step 0 and Step 1.
The missing step is:
To determine the minimum number of times you must throw two fair eight-sided dice to ensure that the same sum is rolled twice, we can use the pigeonhole principle. The pigeonhole principle states that if n items are put into m containers, with n > m, then at least one container must contain more than one item.
Missing Step 2:
The missing step should be placed between Step 1 and Step 2.
The missing step is:
Therefore, the possible sums when rolling two eight-sided dice are the integers from 2 to 16, inclusive. This gives us a total of 16 - 2 + 1 = 15 possible sums.
Qwen2.5-7B-Instruct Missing Step 1:
The missing step should be placed between Step 0 and Step 1.
The missing step is:
The possible sums when rolling two eight-sided dice range from 2 to 16, which gives us 15 different possible sums.
Qwen2.5-72B-Instruct Missing Step 1:
The missing step should be placed between Step 1 and Step 2.
The missing step is:
The possible sums when rolling two eight-sided dice are 2, 3, 4, …, 16. There are 15 possible sums in total.
Missing Step 2:
The missing step should be placed between Step 2 and Step 3.
The missing step is:
According to the pigeonhole principle, if we roll the dice 16 times, there are 15 possible sums, so at least one of these sums must appear at least twice.
Gemini-2.5-PRO Missing Step 1:
The missing step should be placed between Step 1 and Step 2.
The missing step is:
Now, we need to determine the total number of distinct sums possible. The sums are all the integers from 2 to 16, inclusive. These distinct sums will be our "pigeonholes."
To count the number of possible sums:
Number of distinct sums=(Maximum Sum-Minimum Sum)+1
Number of distinct sums=(16-2)+1=14+1=15
So, there are 15 unique possible sums (the pigeonholes) when rolling two eight-sided dice.
GPT-4o Missing Step 1:
The missing step should be placed between Step 1 and Step 2.
The missing step is:
We need to count how many distinct sums can result from rolling two eight-sided dice. The possible sums range from 2 to 16, which gives:
16 - 2 + 1 = 15 distinct possible sums.
As can be seen in the case study:
  • Although a planning-style bridge between the question and step 1 was missing, GPT-4o, Gemini-2.5-PRO, and Qwen2.5-72B-Instruct all identified that "15" lacked a calculation explanation and bridged the gap in the appropriate position.
  • Qwen2.5-72B-Instruct generated a redundant step (almost identical to step 2). This phenomenon is quite common and aligns with its 33.73% redundancy rate shown in Table 3.
  • The consistency of the content and position bridged by Qwen2.5-7B-Instruct is not good.
Appendix J CoT Score
J.1 Score Template
Since scoring a single response with DeepSeek-R1 can result in high variance, we instead compare the original CoT and its bridged version for the same question side-by-side.
DeepSeek-R1 Score Template You are a math reasoning expert. Please evaluate the coherence of the given two Chain of Thoughts (CoTs) based on the provided question.
@@ 5. Conclusion Support: Does it provide a clear summary or verification to support the final conclusion?
Table 10: DeepSeek-R1 scoring results for the original CoT and the bridged CoT.
Appendix K PRM Score